Walter has $7.55 in quarters and dimes. If he has a total of 47 coins, how many are quarters?
Answer by crabbyoldgamer(13)

q + d = 47

.25q + .1d = 7.55


q = 47-d

.25(47-d) + .1d = 7.55

11.75 - .25d + .1d = 7.55

4.2 = .15d

28=d


q = 47-d = 47-28 = 19